### v1

# Optimized quantum sensor networks for ultralight dark matter detection

Abstract
Dark matter (DM) remains one of the most compelling unresolved problems in fundamental physics, motivating the search for new detection approaches. We propose a network-based quantum sensor architecture to enhance sensitivity to ultralight DM fields. Each node in the network is a superconducting qubit, interconnected via controlled-Z gates in symmetric topologies such as line, ring, star, and fully connected graphs. We investigate four- and nine-qubit systems, optimizing both state preparation and measurement using a variational quantum metrology framework. This approach minimizes the quantum and classical Cramér-Rao bounds to identify optimal configurations. Bayesian inference is employed to extract the DM-induced phase shift from measurement outcomes. Our results show that optimized network configurations significantly outperform conventional GHZ-based protocols while maintaining shallow circuit depths compatible with noisy intermediate-scale quantum hardware. Sensitivity remains robust under local dephasing noise. These findings highlight the importance of network structure in quantum sensing and point toward scalable strategies for quantum-enhanced DM detection.
Introduction. Dark matter (DM), comprising about 27% of the universe’s mass-energy content, remains one of the most profound mysteries in modern physics [1, 2]. It is essential for explaining the gravitational binding of galaxies and the formation of large-scale cosmic structures [3, 4]. The concept originated from mass inconsistencies observed in galaxy clusters by Zwicky [5], and was later supported by galactic rotation curves measured by Rubin [6]. Although many experiments target potential DM candidates such as WIMPs, axions, and sterile neutrinos, none have been confirmed [7, 8, 9, 10, 11, 12, 13]. The detection of weakly interacting or light DM candidates, such as axions or hidden photons, remains challenging due to their extremely weak interactions with standard matter [14]. Identifying the fundamental nature of DM remains a central objective in both astrophysics and particle physics.
Recent advances in quantum technologies have opened new avenues for DM detection. Superconducting qubits, in particular, offer exceptional sensitivity to weak external perturbations due to their tunability and long coherence times [15, 16]. It has been proposed that individual qubits can detect ultralight DM through shifts in electromagnetic fields or spin orientations [17], while entangling multiple qubits can enhance the sensitivity [18, 19]. Techniques from quantum metrology, which exploit quantum correlations to surpass classical measurement limits, have been applied to analyze the detection precision [20].
Quantum sensor networks, formed by interconnected sensors, provide advanced capabilities for quantum sensing [21, 22]. By distributing quantum correlations across multiple nodes, such networks can improve both sensitivity and robustness, as demonstrated in star-shaped architectures [23, 24, 25]. However, network topology is crucial, as random or poorly structured configurations may significantly limit sensing performance. Network optimization is crucial for maximizing quantum sensing performance.
In this work, we propose a network-based quantum sensor to enhance DM detection. Each node in the network is a superconducting qubit, connected via controlled-Z (CZ) gates. We explore symmetric network topologies, including line, ring, star, and fully connected graphs, to evaluate their impact on sensing performance. We analyze two representative systems comprising four and nine qubits, respectively.
To optimize sensing performance, we employ a variational quantum metrology (VQM) approach [25, 26, 27], minimizing both the quantum Cramér-Rao bound (QB) and classical Cramér-Rao bound (CB) by variationally updating the initial quantum states and measurement settings. For estimation, we implement a Bayesian inference protocol to extract the DM–induced phase from measurement outcomes.
Our results demonstrate that optimized network-based sensors outperform standard GHZ-based protocols in sensitivity. Optimal configurations achieve high performance with shallow circuits, making them suitable for noisy intermediate-scale quantum (NISQ) devices. The precision improves with increasing qubit number, consistent with theoretical expectations for quantum-enhanced sensing. It also remains robust against dephasing noise.
Beyond DM detection, the optimization methods and network architectures developed here are broadly applicable to a range of quantum sensing tasks, including gravitational wave detection, precision spectroscopy, and magnetic field sensing [28]. By integrating quantum computing, metrology, and fundamental physics, this approach provides a practical framework for addressing some of the most challenging problems in science.
DM detection via superconducting qubits. Recent studies have used superconducting transmon qubits as quantum sensors for detecting ultralight DM [16, 17, 18, 19]. In these schemes, a coherently oscillating DM field induces an effective electric field, $\bar{E}^{(\text{eff})}$, which drives Rabi-like transitions between the ground and excited states of a transmon qubit. The interaction is described by the effective Hamiltonian
$\mathcal{H}=\omega|e\rangle\langle e|-2\eta\cos\big{(}m_{\text{DM}}t-\alpha\big{)}\big{(}|e\rangle\langle g|+|g\rangle\langle e|\big{)},$ (1)
where $\omega$ is the qubit energy splitting, $m_{\text{DM}}$ is the DM mass, $\alpha$ is the DM phase offset, and $|g\rangle$, $|e\rangle$ denote the ground and excited states, respectively. The coupling strength $\eta$ is given by
$\eta\equiv\frac{\sqrt{C\omega}}{2\sqrt{2}}\bar{E}^{(\text{eff})}d\cos\Theta,$ (2)
where $C$ and $d$ are the transmon’s capacitance and plate separation, and $\Theta$ is the angle between $\vec{E}^{(\text{eff})}$ and the normal vector of the capacitor plates. In simulations, $\cos^{2}\Theta\approx 1/3$ to account for angular fluctuations within the DM coherence time [18, 19].
The evolution of the qubit state is governed by the time-dependent Schrödinger equation
$i\frac{d}{dt}|\Psi(t)\rangle=\mathcal{H}|\Psi(t)\rangle,$ (3)
where $|\Psi(t)\rangle=\psi_{g}(t)|g\rangle+\psi_{e}(t)|e\rangle$ denotes the qubit wavefunction expressed in the ground and excited state basis. Within the coherence time $\tau\simeq\min(\tau_{\text{DM}},\tau_{q})$, where $\tau_{\text{DM}}$ and $\tau_{q}$ are the coherence times of the DM field and the qubit, respectively, the evolution is described by a unitary operator $U_{\text{DM}}(\tau)$
$\begin{pmatrix}\psi_{g}(\tau)\\ \psi_{e}(\tau)\end{pmatrix}=U_{\text{DM}}(\tau)\begin{pmatrix}\psi_{g}(0)\\ \psi_{e}(0)\end{pmatrix}.$ (4)
In the resonant regime, $\omega=m_{\text{DM}}$, the interaction simplifies under the rotating wave approximation, and the evolution operator takes the form
$U_{\text{DM}}(\tau)=\begin{pmatrix}\cos\delta&ie^{-i\alpha}\sin\delta\\ ie^{i\alpha}\sin\delta&\cos\delta\end{pmatrix},\quad\delta\equiv\eta\tau,$ (5)
with eigenstates $|\psi_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|g\rangle\pm e^{i\alpha}|e\rangle\right).$
In the individual-qubit readout scheme, each transmon functions as an independent sensor, and the excitation probability induced by the DM field accumulates linearly with the number of qubits $p_{g\rightarrow e}(\tau)\simeq N\delta^{2}$, for $N\delta^{2}\ll 1,$ where $N$ is the number of qubits.
To surpass this shot-noise limit, quantum-enhanced sensing protocols based on entangled states such as GHZ states have been proposed [18, 19]. At the initial time $t_{1}$, a GHZ state is prepared
$|\Psi(t_{1})\rangle=\frac{1}{\sqrt{2}}\Big{(}|g\rangle^{\otimes N}+|e\rangle^{\otimes N}\Big{)}.$ (6)
During interacting with the DM field over a time interval $\tau=t_{2}-t_{1}$, each qubit acquires a phase shift $\delta$, resulting in the global state
$|\Psi(t_{2})\rangle=\frac{1}{\sqrt{2}}\Big{(}e^{iN\delta}|+\rangle^{\otimes N}+e^{-iN\delta}|-\rangle^{\otimes N}\Big{)},$ (7)
where |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|g\rangle\pm|e\rangle)$.
After the interaction, a reversed GHZ operation is applied to coherently map the accumulated phase information onto a single qubit
$|\Psi(t_{f})\rangle=\Big{(}\cos(N\delta)\,|g\rangle+i\sin(N\delta)\,|e\rangle\Big{)}\otimes|+\rangle^{\otimes(N-1)}.$ (8)
Here, we assume $t_{f}=t_{2}$, neglecting further evolution beyond the sensing interval. The entangled state enables phase accumulation at a rate proportional to $N^{2}\delta^{2}$, yielding sensitivity that scales quadratically with $N$, in contrast to the linear scaling of individual-qubit strategies. This quadratic enhancement highlights a central advantage of entangled probes in quantum metrology, enabling parameter estimation beyond the standard quantum limit.
However, the optimal state for enhancing metrological precision is a superposition of the minimum and maximum eigenstates of the unitary evolution operator $U_{\rm DM}$, taking the form of a GHZ-like state [29]
$|\Psi(t_{1})\rangle=\frac{1}{\sqrt{2}}\Big{(}|\psi_{+}\rangle^{\otimes N}+|\psi_{-}\rangle^{\otimes N}\Big{)}.$ (9)
In contrast, the conventional GHZ state in Eq. (6) does not yield improved precision, as shown by the QB in Fig. 2(a). To overcome this limitation, we employ a network-based quantum sensor architecture to rigorously evaluate the achievable sensitivity. We use a VQM approach to optimize performance beyond what is attainable with GHZ states alone.
Figure 1: (a) A sensor-network architecture for detecting DM, which composes a 2D array of superconducting qubits connected via controlled-Z (CZ) gates to form a tunable interaction network. (b) Network configurations for $N=4$ qubits. (c) Network configurations for $N=9$ qubits. (d) Schematic of a VQM framework. The ansatze $U(\bm{\theta})$ (quantum state preparation) and $M(\bm{\mu})$ (measurement) are drawn from the structures in (b) and (c). The unitary $V(\delta,\alpha)$ encodes the external DM-induced evolution. The cost function $\mathcal{C}(\bm{\theta},\bm{\mu})$ guides the optimization based on QB and CB. The reconstruction is given through Bayesian estimation. (e) Example variational circuit structure for S4 configuration, showing the placement of single-qubit rotations and CZ gates.
Quantum-enhanced metrology viewpoint. The precision in estimating $\delta$ is fundamentally bounded by Cramér-Rao bounds
$\nu\,\text{Var}(\delta)\geq\frac{1}{F(\delta)}\geq\frac{1}{Q(\delta)},$ (10)
where $\nu$ is the number of independent measurements, $F(\delta)$ is the classical Fisher information (CFI), and $Q(\delta)$ is the quantum Fisher information (QFI). These bounds quantify the ultimate limits on parameter estimation precision, with the QFI setting the best achievable sensitivity for any measurement strategy.
CFI depends on the measurement outcomes through the conditional probability distribution $p(m|\delta)$, and is given by
$F(\delta)=\sum_{m}\frac{1}{p(m|\delta)}\left(\frac{\partial p(m|\delta)}{\partial\delta}\right)^{2}.$ (11)
This expression quantifies the information about $\delta$ obtainable from a specific measurement strategy.
In contrast, QFI characterizes the maximum information attainable over all possible quantum measurements. It depends solely on the quantum state $\rho_{\delta}$ and is defined as
$Q(\delta)=\text{Tr}\left[\rho_{\delta}L_{\delta}^{2}\right],$ (12)
where $\rho_{\delta}=|\Psi(t_{2})\rangle\langle\Psi(t_{2})|$ is the state after sensing and $L_{\delta}$ is the symmetric logarithmic derivative (SLD), determined by $2\partial\rho_{\delta}/\partial\delta=L_{\delta}\rho_{\delta}+\rho_{\delta}L_{\delta}$. A larger QFI indicates greater sensitivity, enabling more precise parameter estimation.
For separable (non-entangled) states, the QFI scales linearly with the number of qubits, $Q(\delta)\propto N,$ corresponding to the shot-noise limit, which represents the fundamental precision bound for classical strategies.
In contrast, when using an entangled GHZ state, the QFI scales quadratically with the number of qubits, $Q(\delta)\propto 4N^{2},$ demonstrating Heisenberg-limited sensitivity. However, as noted above, standard GHZ states do not always yield optimal precision. This motivates further exploration of adaptive protocols and optimal measurement strategies, as discussed below.
Figure 2: (a) Comparison of QBs achieved by different ansatz configurations for a four-qubit system with $L_{1}=1$. The optimized ansatze, particularly the fully connected F4 architecture, outperform standard choices like the GHZ and excited states. (b) QB performance of F4 configuration as a function of the number of layers $L_{1}$. The best precision is achieved with a shallow circuit ($L_{1}=1$), highlighting the efficiency and practicality of the method for NISQ devices. (c) Robustness of the optimal configuration (F4, $L_{1}=1$) across a wide range of $\delta$ and $\alpha$. The QB remains nearly constant with $\delta$ and exhibits periodic variation with $\alpha$, demonstrating adaptability without the need for re-optimization.
Quantum networks for DM detection. We consider a network-based quantum sensor modeled as a two-dimensional array of superconducting qubits, as illustrated in Fig. 1(a). Each qubit acts as a node, with edges representing physical connections between nodes. Due to hardware constraints, each qubit is limited to a maximum of four connections, consistent with the restricted connectivity in current superconducting architectures. We focus on symmetric network topologies known to enhance sensing precision. Specifically, we examine systems of four and nine qubits, shown in Fig. 1(b) and Fig. 1(c), respectively. For the 4-qubit case, we analyze linear (L4), star (S4), ring (R4), and fully connected (F4) configurations. For the 9-qubit system, we study analogous structures with linear (L9), star (S9), ring-star hybrid (RS9), and fully connected network (F9). These network architectures provide a structured platform for exploring the role of geometry in quantum sensing performance.
Variational quantum metrology. VQM is a hybrid quantum-classical framework for optimizing quantum sensing protocols using parameterized quantum circuits [25, 26, 27]. In this scheme, an ansatz $U(\bm{\theta})$ prepares the quantum state. Following interaction with the target process, such as the DM-induced unitary $U_{\rm DM}$, a second ansatz $M(\bm{\mu})$ is applied to optimize the measurement basis. These ansatze are governed by trainable parameter sets $\bm{\theta}$ and $\bm{\mu}$, respectively.
The optimization loop employs an optimizer to update $\bm{\theta}$ and $\bm{\mu}$ by minimizing (or maximizing) a cost function, which may be the QB, CB, or another task-specific metric. This hybrid framework enables adaptive learning of optimal sensing strategies and has been experimentally demonstrated on various quantum platforms [30, 31]. It is well suited for near-term devices that lack full error correction.
In this work, we implement the VQM scheme illustrated in Fig. 1(d), where the ansatze $U(\bm{\theta})$ and $M(\bm{\mu})$ are structured according to the network topologies shown in Figs. 1(b) and 1(c). These ansatze consist of rotation blocks $U_{\rm rot}=R_{z}(\theta_{z})R_{y}(\theta_{y})R_{x}(\theta_{x})$, inserted between each controlled-Z (CZ) gate, which defines the qubits connectivity [25]. An explicit example of the ansatz $U(\bm{\theta})$ for S4 configuration is provided in Fig. 1(e). We denote $L_{1}$ and $L_{2}$ as the number of layers in $U(\bm{\theta})$ and $M(\bm{\mu})$, respectively.
The sensing process is described by the global unitary operator
$V(\delta,\alpha)=\bigotimes_{i=1}^{N}U_{\rm DM}(\delta,\alpha),$ (13)
which models interactions between each qubit and the external DM field. Here, $U_{\rm DM}(\delta,\alpha)$ is the single-qubit evolution operator defined in Eq. (5).
Optimization and estimation. We begin by optimizing the parameters $\bm{\theta}$ by minimizing a cost function defined by the QB $\mathcal{C}(\bm{\theta})=Q^{-1}(\delta),$ yielding the optimal configuration $\bm{\theta}^{*}=\arg\min_{\{\bm{\theta}\}}\mathcal{C}(\bm{\theta})$. Once $\bm{\theta}^{*}$ is fixed, we optimize the measurement ansatz $M(\bm{\mu})$ using a similar procedure, with the cost function based on the CB, $\mathcal{C}(\bm{\mu})=F^{-1}(\delta).$ To update parameters, we use Adam optimizer.
After optimizing the network state and measurement configuration, we apply Bayesian estimation to infer the unknown parameter $\delta$ from the measurement outcomes. Let $m\in\{00\cdots 0,\ldots,11\cdots 1\}$ denote an outcome in the computational basis, and $p(\delta)$ the prior probability distribution representing initial knowledge of $\delta$. Here, we use the probability density function (pdf). The likelihood function $p(m|\delta)$ is given by the probability of observing outcome $m$ for a given $\delta$, computed from the final quantum state.
Using Bayes’ theorem, the posterior distribution $p(\delta|m)$ is given by
$p(\delta|m)=\frac{p(m|\delta)p(\delta)}{p(m)},$ (14)
[…]
The posterior distribution $p(\delta|m)$ encodes all information about $\delta$ after the measurement. The estimate $\bar{\delta}$ is given by the posterior mean, and the precision is quantified by the variance $\mathrm{Var}(\delta)$.
Numerical results. For VQM optimization, we fix $\delta=0.05$ and $\alpha=0$. The parameters $\bm{\theta}$ and $\bm{\mu}$ are initialized randomly. Measurement outcomes are simulated using the Primitives sampler from Qiskit.
Quantum bound. We evaluate the performance of various ansatz configurations for a four-qubit system ($N=4$) and present the results in Fig. 2(a) for $L_{1}=1$. The GHZ and excited state (E) do not achieve QBs comparable to those of the optimized configurations. As noted earlier, the GHZ state is suboptimal under evolution by $U_{\rm DM}$ because it is not aligned with the eigenstates associated with the eigenvalues of the unitary. Consequently, its performance is comparable to that of the excited state.
By contrast, the optimized ansatze, particularly those with more connections such as L4, R4, S4, and especially F4, outperform the others. The best performance of F4 is attributed to its higher connectivity and larger number of trainable parameters, which significantly enhance quantum state correlations and contribute to greater circuit expressiveness.
In Fig. 2(b), we explore the QB of F4 for different layers $L_{1}$. Remarkably, the best performance is achieved with just a single layer ($L_{1}=1$). This result is promising, as it implies that high precision can be obtained using shallow circuits, making the method both efficient and suitable for NISQ devices.
With the optimal configuration fixed at F4, $L_{1}=1$ and using the optimized parameters $\bm{\theta}$ found earlier, we next examine the QB as a function of the parameters $\delta$ and $\alpha$, as shown in Fig. 2(c). We observe that for a fixed $\alpha$, the QB remains constant over a wide range of $\delta$, from $10^{-3}$ to $10^{0}$. This robustness implies that re-optimization is not required for each new value of $\delta$, offering significant practical advantages for sensing the DM field. When varying $\alpha$, the QB shows periodic behavior, peaking at $\alpha=\pi/2$ and returning to its original value at $\alpha=0$ and $\pi$. Since the DM phase offset $\alpha$ is typically under experimental control, this does not negatively affect the practical performance.
These findings highlight the potential of our network-based approach to generate high-performance quantum resources for metrology and sensing. Next, we optimize CB.
Figure 3: (a) CB as a function of the number of layers $L_{2}$ in the measurement circuit $M(\bm{\mu})$. Increasing $L_{2}$ improves the CB, bringing it closer to the QB, until an optimal point is reached. Beyond this, further increasing $L_{2}$ causes performance degradation. (b) CB versus different values of $\delta$ and $\alpha$. The optimized measurement configuration maintains a nearly constant and minimal CB across a wide range of $\delta$, except for relatively large $\delta$ where CB slightly increases. A periodic dependence on $\alpha$ is observed, consistent with the behavior seen in the QBs.
[…]
Figure 3(a) shows that with a shallow measurement circuit ($L_{2}=1$), the CB is larger than the QB, indicating poor measurement performance. As $L_{2}$ increases, the CB approaches the QB, reflecting improved accuracy. However, beyond an optimal depth, further increasing $L_{2}$ leads to a rise in the CB, reducing precision. This non-monotonic behavior suggests that deeper circuits may suffer from limited expressibility, overparameterization, or barren plateaus, where the optimization landscape becomes flat. The gap between CB and QB arises from the inability of the measurement ansatz to fully approximate the optimal POVM associated with the QFI.
In Fig. 3(b), we examine how the CB varies with different values of $\delta$ and $\alpha$. The results show that our optimized measurement configuration yields a constant, minimal CB over a wide range of $\delta$, similar to the behavior observed for the QB in Fig. 2(c). However, when $\delta$ becomes relatively large (on the order of $10^{0}$), the CB begins to increase. This rise does not significantly impact our conclusions, as the physically relevant regime for the DM field corresponds to small $\delta$ values.
Regarding $\alpha$, the phase offset of the DM interaction, we observe periodic variations in the CB, consistent with the behavior of the QB. This periodicity indicates that estimation performance depends on $\alpha$, but the optimal configuration remains robust across its range.
Bayesian estimation. Figure 4 presents the results of the Bayesian estimation procedure. In Fig. 4(a), we plot the posterior probability distribution as a function of $\delta$ for different numbers of measurements $\nu$. As $\nu$ increases, the peak of the posterior distribution shifts closer to the true value of $\delta$, demonstrating improved estimation accuracy.
Figure 4(b) shows the bias $\bar{\delta}-\delta_{\rm true}$ versus $\nu$. The bias gradually approaches zero as $\nu$ increases, demonstrating that the estimator becomes asymptotically unbiased with more measurement data.
Figure 4: (a) Posterior probability distribution versus the phase $\delta$ for different number of measurements $\nu$. As $\nu$ increases, the posterior peak moves closer to the true value. (b) Bias $\bar{\delta}-\delta_{\rm true}$ as a function of $\nu$, showing convergence to zero with increasing $\nu$. (c) Variance Var($\delta$) versus $\nu$, compared with the ${\rm CB}/\nu$ and ${\rm QB}/\nu$. The results demonstrate good agreement, confirming the optimal scaling behavior.
In Fig. 4(c), we plot Var($\delta$) against $\nu$ and compare it with the theoretical predictions given by the QB and CB, both scaled as $1/\nu$. The results show excellent agreement: the variance closely follows the trends of ${\rm QB}/\nu$ and ${\rm CB}/\nu$, confirming that our method achieves the optimal scaling expected from quantum estimation theory. For larger $\nu$, the estimator becomes unbiased, and the bounds in Eq. (10) hold.
For $N=9$. We evaluate the QB and present the results in Fig. 5. Using the same procedure as in Fig. 2(a) for the four-qubit system, we examine multiple configurations shown in Fig. 1(c). Among these, the fully connected architecture F9 delivers the highest QB performance.
When compared to the four-qubit case, the QB for $N=9$ is significantly improved. This improvement is expected because increasing the number of qubits generally enhances the achievable precision in quantum sensing, as predicted by quantum metrology theory.
In Fig. 5(b), we further investigate the QB as a function of $L_{1}$. Different from four-qubit case, here, the best performance is obtained when ($L_{1}=2$). This result suggests that even for larger systems, a shallow and well-optimized circuit is sufficient to achieve high-precision estimation, which is advantageous for practical implementation on near-term quantum devices.
Figure 5(c) shows how the CB varies with $L_{2}$. As $L_{2}$ increases, the CB improves and approaches the QB. However, it does not reach the QB due to limitations in the measurement strategy. In this analysis, we evaluate up to $L_{2}=3$. Remarkably, the precision in this case is better than that of the four-qubit case.
Figure 5: QBs for a nine-qubit system ($N=9$). (a) Comparison of various ansatze, following the designs shown in Fig. 1(c). Among them, F9 achieves the best QB. Compared to the four-qubit case, the QB is improved, consistent with the expectation that precision increases with the number of qubits. (b) Dependence of the QB on the number of layers $L_{1}$ for F9 configuration. The best performance is achieved with a shallow circuit at $L_{1}=2$, highlighting the efficiency and practicality of the approach for larger quantum networks. (c) CB as a function of the number of layers $L_{2}$. Increasing $L_{2}$ improves the CB, bringing it closer to the QB.
Noisy effect. We analyze the efficiency of a quantum sensor network under local dephasing noise. Each sensor node (qubit) undergoes dephasing described by the Kraus operators $K_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda}\end{pmatrix},K_{1}=\begin{pmatrix}0&0\\ 0&\sqrt{\lambda}\end{pmatrix}$, where $\lambda\in[0,1]$ characterizes the dephasing strength. The resulting channel acts as
$\mathcal{E}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}.$ (15)
The numerical results for the QB under varying dephasing strength $\lambda$ are shown in Fig. 6 for two configurations, F4 (a) and F9 (b). No re-optimization is performed for each noise level; instead, the optimal configuration setup at $\lambda=0$ is used throughout. Remarkably, the QBs slightly increase even as the noise strength rises to $\lambda=0.9$. Specifically, the difference between the QB at $\lambda=0.9$ and the noiseless case is $\Delta=1.68\times 10^{-5}$ for F4 and $\Delta=3.73\times 10^{-5}$ for F9. This demonstrates that the optimal sensor networks remain robust under dephasing, with only minimal degradation in sensitivity even without re-optimization.
Figure 6: QBs under dephasing noise for (a) F4 and (b) F9 configurations. QBs exhibit a slight increase with increasing dephasing strength $\lambda$, indicating a modest reduction in sensitivity. However, even at high noise levels ($\lambda=0.9$), the increase remains small: $\Delta=1.68\times 10^{-5}$ for (a) and $\Delta=3.73\times 10^{-5}$ for (b).
Conclusion. We proposed and demonstrated a network-based quantum sensing strategy for enhancing DM detection. By structuring superconducting qubits into networks with specific symmetries and optimizing their connections using VQM, we achieved substantial improvements in sensing precision compared to conventional GHZ-based protocols. Both the QB and CB were systematically minimized, and Bayesian estimation further validated the enhanced sensitivity of the optimized networks. Importantly, the optimal configurations maintained shallow circuit depths, making them feasible for implementation on NISQ devices. Our findings emphasize the significant impact of network design on quantum metrology performance and suggest that well-structured quantum networks can serve as powerful platforms not only for DM searches but also for a wide range of high-precision sensing applications. Future works may extend these results by exploring larger networks, adaptive strategies, and integration with error mitigation techniques to further push the limits of quantum-enhanced detection.
Acknowledgments. This paper is supported by JSPS KAKENHI Grant Number 23K13025.
Data availability. No data were created or analyzed in this study.